a+b=3
-b+c=3
a+2c=10

Answer by rahman(247)

a+b=3 ...............(1)
-b+c=3 ...............(2)
a+2c=10 ...............(3)
Use elimination method for (1) and (2):
a+b=3
-b+c=3
------ +
a+c=6 .............(4)
Use elimination method for (3) and (4):
a+2c=10
a+c=6
------ -
c=4
Substitute c=4 to (4):
a+c=6
a+4=6
a=6-4
a=2
Substitute a=2 to (1):
a+b=3
2+b=3
b=3-2=1
Thus, a=2, b=1, and c=4.
OK